OpenStudy (anonymous):

When Justin goes to work, he drives at an average speed of 55 miles per hour. It takes about 1 hour and 30 minutes for Justin to arrive at work. His car travels about 30 miles per gallon of gas. if gas costs $3.45 per gallon, how much money does Justin spend to travel each mile to work?

OpenStudy (superdavesuper):

u need to find the distance traveled first....distance = speed x time so how far he has traveled?

OpenStudy (anonymous):

(55)(90) = 4950 miles

OpenStudy (superdavesuper):

close but not quite - the speed 55 is in miles per hour so the time needs to be in hour as well....

OpenStudy (anonymous):

(55)(1.5) = 82.5

OpenStudy (superdavesuper):

so u need to calculate how many gallons of gas are used by this - "30 miles per gallon of gas"

OpenStudy (anonymous):

82.5 miles / 30 miles per gal = 2.75 * 3.45 = 9.49 spent traveling each mile to work

OpenStudy (whpalmer4):

it's best to keep the units to prevent these sorts of errors: 55 miles/hr * 90 minutes = 4950 miles * minutes/hr — funky unit is sign that you haven't done this right 55 miles/hr * 90 minutes * (1 hr/60 minutes) = = 4950/60 * miles/hr * minutes * hr/minutes = 82.5 miles — a much more sensible unit!

OpenStudy (superdavesuper):

then u multiple the gallons of gas by price "gas costs $3.45 per gallon"....yup yup hehee u are ahead of me!

OpenStudy (superdavesuper):

$9.49 is the right ans good job! :)

OpenStudy (whpalmer4):

again, write the work with the units and do the cancellation carefully! 82.5 miles * 1 gallon / 30 miles = 82.5/30 gallons = 2.75 gallons 2.75 gallons * $3.45 /gallon = $9.49 for the trip to work but the problem asks for cost PER MILE

OpenStudy (whpalmer4):

$9.49/82.5 miles =

OpenStudy (superdavesuper):

Ooops...@whpalmer4 is right....sorry i thought it was asking for the total cost :(

OpenStudy (whpalmer4):

this is why I always reread the problem before checking my answer. occasionally discover that I've solved the wrong problem :-)

OpenStudy (superdavesuper):

come back @pmcafee u need to do one more thing :)

OpenStudy (anonymous):

.12 spent to travel EACH MILE to work.

OpenStudy (anonymous):

thank you so much.

OpenStudy (superdavesuper):

welcome :)

OpenStudy (superdavesuper):

sorry about my earlier mistake!

OpenStudy (anonymous):

I posted two other questions.